OpenStudy (anonymous):

Find a rational function f: R ->R with range f(R) = [0,1]. (Thus f(x) = p(x)/q(x) for all x for suitable polynomials P and Q where Q has no real roots.

OpenStudy (anonymous):

try using a square root somewhere in the answer.

OpenStudy (anonymous):

or... an exponentiation!

OpenStudy (anonymous):

i have literally no idea to start, can you help me start off first please?

OpenStudy (anonymous):

thanks, I have to head to a lecture now but will look at it later!

OpenStudy (anonymous):

Still doesn't work... :( oh well let Q be defined as \[(\sqrt{x})^2 - 1\]

OpenStudy (anonymous):

hmm that's still not right

OpenStudy (anonymous):

\[Q(x) =\sqrt{1-x}\]

OpenStudy (anonymous):

almost there

OpenStudy (anonymous):

alright I got it :\[Q(x) = \sqrt{1-x^2}\]

OpenStudy (anonymous):

P(X) can be anything,.

OpenStudy (anonymous):

Why don't you take \(q(x)=x^2+a^2\), for any real number a?

OpenStudy (zarkon):

\[f(x)=\frac{2x^2}{x^4+1}\]

OpenStudy (anonymous):

Zarkon, could you have a look at the problem right below this one?

OpenStudy (anonymous):

thanks guys. Q(x) = \[\sqrt{1-x ^{2}}\] and P can equal anything for example \[x^{3}\] ?